# Inference of time delay in stochastic systems

### v1

@@ A.1 Numerical solution of the overdamped delay Langevin equation
where $N_{\tau}=\tau/\Delta t$ is the number of time steps within one delay time $\tau$, for the discretization time step $\Delta t$. $\eta_{n}$ is a uncorrelated Gaussian random number with zero mean and unit variance, i.e., $\langle\eta_{n}\eta_{m}\rangle=0$ for $n\neq m$. Equation (19) is solved provided an initial history for the time interval $[-\tau,0]$ via the discretized history function $\Phi(t_{n})=\Phi_{t_{n}}$, that is, the trajectory in the discretized time interval $[-\tau,0]$, has to be known. For all stochastic simulations in this study, we used the trajectories of free Brownian particles as the initial history. Furthermore, we used time steps of $\Delta t=10^{-4}\tau_{B}$ (much smaller than the smallest delay time considered). to ensure sufficient numerical precision in all the cases considered.
A.2 The Differencing method
@@
Figure A1 presents exemplary data for systems with attractive and repulsive harmonic time-delayed feedback which both behave diffusively, especially in the long time limit. The cyan and red data correspond to PSD computed by directly applying the Fourier transform to the nonstationary positional data.
Figure A1: Comparison of position based and differencing based computation of $\mathrm{PSD}$ and $\omega^{2}\mathrm{PSD}$ for a system with linear feedback force [Eq. (2)]. Panels (a) and (c): Marginally stable system with attractive harmonic feedback; panels (b) and (d) are for repulsive harmonic feedback below the deterministic threshold, i.e., $|k_{a}|=|k_{b}|<\gamma/\tau$.
We observe a shift of compared to the analytical expectation that is hardly visible in the PSD but becomes apparent in the $\omega^{2}$. While for averaging over a small number of noise realizations this shift is lost in noise, when averaging over a larger number of noise realizations the shift becomes obvious.
The reason behind this behavior can be understood when considering that we are taking the Fourier transform of a (within the observation period) not strictly ergodic and certainly non-stationary quantity. More specifically, the time series is non-stationary since the variance of the position grows over time. Changing the temporal resolution or the length of the trajectory only has a small effect on the observed shift. The observed shift can intuitively be understood as follows: In a stationary system (assuming zero mean and constant variance), the position will (more or less) fluctuate around zero, given that the observation window is long enough. However, if the process is not stationary in the sense that the variance grows as a function of time, in some realizations, the position will be positive (or negative) for a significant part of the observation time. When taking the Fourier transform of such a trajectory, it will be somewhat biased. If we took the average over noise realizations now, we could expect this bias to average out over different realizations. However, following the definition of the PSD (8) we compute the square modulus of the Fourier transform of the position before applying the noise average. For a stationary process, this does not matter, but in our case, the bias can now not be averaged out but rather accumulates and depends on the variance of the process and how much of the observation period will be uninterruptedly spent on one side of the mean position. We therefore have to approach the problem differently and make the time series stationary before even computing the Fourier transform in order to obtain results that match our analytical expectation.
@@
where the operator $\mathcal{F}$ represents the Fourier transform. Intuitively, this looks like the Fourier transform over a temporal derivative (which we already exploited in the analytical considerations) or essentially (for small $\Delta t$) the Fourier transform of the right-hand side of the discretized Langevin equation. Applying this method, we can obtain a numerical result that matches exactly the analytical expectation, as shown by the dark blue curves in Fig. A1.
A.3 Extraction of frequencies corresponding to minima and maxima of the PSD
In the following, we describe a procedure that allows us to extract the delay time from simulation data. It is based on the analytical results obtained for the linear system, but can also be applied to systems that depend on $x(t)$ and/or $x(t-\tau)$ in a nonlinear fashion. This procedure relies on the spacing of the positions of maxima and minima in the quantity $\omega^{2}\mathrm{PSD}$, and is applicable as long as the signal-to-noise ratio in the system allows for reliable determination and extraction of the position of these minima and maxima. As such, the larger the number of available noise realizations, the better the inference results. For example, Fig. A1 illustrates the level of noise that has to be dealt with in this kind of system in the case of a marginally stable time delay system. In order to obtain a good polynomial fit over multiple peaks, we need to average over a sufficient number of noise realizations; for the figures in the main text [Figs. 6 400 noise realizations were used.
While, in principle, the positions of the minima can be obtained manually, it is more reliable and convenient to use a spline interpolation for a subsequent semi-automatic extraction of the positions of minima and maxima. It is semi-automatic in the sense that one has to determine for how many minima and maxima a reliable fit can be obtained. Specifically, we use quartic splines for fitting the numerical data, which allows us to extract the roots of its derivative automatically. As the data are noisy and somewhat irregular, a special care has to be taken when adjusting the smoothing parameter in the spline interpolation, such that the interpolation faithfully represents the location of the minima and maxima in the numerical data. Furthermore, limiting the range of the interpolation can significantly improve the results since at low frequencies additional oscillations with different oscillation period are observed in some cases and at high frequencies the oscillations are lost in noise.
@@ A.3 Extraction of frequencies corresponding to minima and maxima of the PSD
    Compute the roots $\omega_{n}$ of the derivative of the polynomial corresponding to the minima and maxima in $\omega^{2}PSD$ (built-in function of spline interpolation in scipy).
@@
$2\pi/\tau_{\mathrm{i}}=\left(\sum_{n=1}^{N_{\mathrm{ext}}-2}\omega_{n+2}-\omega_{n}\right)/(N_{\mathrm{ext}}-2)$ (21)
where $N_{\mathrm{ext}}$ is the number of minima and maxima that can be reliably identified, and we averaged the distances between all available pairs of minima and maxima.
A.4 Deep learning: data preparation
@@
The data are prepared in this way for training the neural network as well as for inference using the trained network.
A.5 Deep learning: network architecture and hyperparameter tuning
For the classification of trajectory data we use a convolutional deep neural network (see e.g., [43, 44]) implemented using keras [81] based on tensorflow [82]. The neural network (see Fig. 9 for a simplified schematic depiction) consists of an input layer (the number of nodes equals the number of relevant time steps), hidden layers, and an output layer (the number of nodes equals the number of classes, i.e., delay times the network is trained on). For hidden layers, considerations such as complexity, regularization, and available resources play a crucial role. Here, we focus on a few convolutional layers and pooling layers to limit complexity. We use a hyperparameter tuning scheme in order to achieve the desired accuracy for classification by the network. This means that whether pooling is applied at all in a pooling layer and over how many nodes, filter and kernel size in convolutional layers, the optimization method and the activation function are chosen automatically. Before the final hidden layer, the network is flattened. The final hidden layer is then a densely connected layer followed by the densely connected output layer where the number of nodes is equal to the number of different delay times on which the model is trained and where a softmax activation function is used to obtain a classification in terms of the probability that the classified trajectory is based on a certain delay time. The network is then constructed as shown in Table 2.
@@ A.5 Deep learning: network architecture and hyperparameter tuning
Table 2: Network structure by layers; hyper-parameters, i.e., whether pooling is applied at all in a pooling layer and over how many nodes, filter and kernel size in convolutional layers, the optimization method and the activation function are optimized by hyperparameter tuning. For the input data preparation, see Appendix A.4.
@@
As the loss function, which is minimized during training, the sparse categorical cross-entropy was used. Figure A2 presents exemplary accuracy [Fig. A2(a)] and loss [Fig. A2(b)] as a function of training epochs for the training of the neural network on several nonlinear feedback forces [Eqs. (12)-(14)] simultaneously. We observe good convergence already at as little as 20 training epochs.
@@
Figure A2: Training the convolutional neural network on nonlinear feedback forces [Eqs. (12)–(14)]: Panel (a) shows the classification accuracy, and panel (b) shows the training loss (computed via sparse categorical crossentropy), both plotted as functions of training epochs for the training and validation datasets.
A.6 Characteristic response of a nonlinear feedback force system
@@
Figure A3 presents exemplary data for the response of systems with nonlinear feedback forces. A complex but still characteristic response is observed.
@@
Figure A3: Characteristic response of a nonlinear system with time delay to an external perturbation (see also Sec.V.1). Panel (a): Response of the $\tanh$ feedback force [Eq.(3)] to a constant external force. Panel (b): The response of a system based on Eq. (14) to the same type of perturbation. Vertical dotted lines indicate the onset and termination of the perturbation, while the vertical dashed line marks the end of the observation window.
Appendix B Analytical details
B.1 Definition of the power spectral density via correlation functions
For context we note that the power spectral density is often defined via the Fourier transform of the corresponding autocorrelation function [86] under the assumption of stationarity of the underlying stochastic process. For one trajectory, the power spectral density is defined as
[…]
×$\hphantom{=}\times x_{\mathcal{T}}(t)dt\bigg{]}e^{-i2\pi ft^{\prime}}dt^{\prime}$ (25)
[…]
with $R_{xx}$ being the autocorrelation function of the position $x(t)$. For e.g., a confined Brownian particle, the mean and variance of the position are constant after some time. Under this condition, the power spectral density can be computed via the Fourier transform of the corresponding positional autocorrelation function as well as via the modulus of the Fourier transform of the position. We fully rely on the latter since the autocorrelation is only of limited use in analytical computations for time delay systems considered here. Furthermore, the Fourier transform of the position of a linear system can be easily computed both analytically and numerically.
B.2 Linear time-delayed feedback - small frequency behavior
where we defined
$\tau_{\pm}\equiv\gamma\frac{1\pm\sqrt{1-k_{a}/k_{b}}}{k_{a}}\ .$ (28)
From the approximation (Eq. 27) we can see that the $\mathrm{PSD}$ depends on $1/\omega^{2}$. For $k_{a,b}>0$, this behavior can be interpreted as a reduced ability of the feedback trap to confine compared to the instantaneous harmonic trap, resulting in a corresponding deviation from the black curve in Fig. 1 at small $\omega$ due to diffusive behavior in the long time limit. Analogously, for $k_{a,b}<0$ the frequency dependency is connected to enhanced diffusion due to repulsive time-delayed feedback
B.3 Linear time-delayed feedback - marginally stable system
B.4 Gaussian time-delayed feedback - analytical details
B.4.1 Expansion of the nonlinear feedback force in the limit of small displacements
For small displacements, that is, small $A$ and $\tau$, the feedback force can be linearized [50], analogously to a small delay approximation [52], resulting in a force of the form of Eq. (2), where $k_{a}=k_{b}=A/b^{2}$.
B.4.2 Expansion of the nonlinear feedback force in the constant velocity state
We can exploit the marginally stable constant velocity state [50] this kind of feedback force can produce and rewrite the equation of motion in terms of position $x(t)$ relative to the position of a particle moving at constant velocity $v_{\infty}$. We can then linearly approximate the feedback force, assuming that the deviation is small. This leaves us with an approximate equation of motion that describes the stochastic dynamics in the ‘constant velocity state’ [50]:
$\gamma\dot{x}\approx v_{\infty}+K\left[x(t)-x(t-\tau)-v_{\infty}\tau\right]+\xi,$ (29)
where we define
$\tilde{x}_{v_{\infty}}=\frac{(\gamma+K\tau)v_{\infty}\delta(\omega)+\tilde{\xi}}{\gamma i\omega-K+Ke^{-i\omega\tau}}.$ (32)
In our simulations, we cannot resolve the zero-frequency contribution, which we therefore neglect in the following. Under this assumption, the Fourier transform of the position would be the same as for the linear feedback force with $k_{a}=k_{b}>0$, which was discussed in the main text. We note that persistent motion cannot be considered a stationary state but rather a non-equilibrium steady state. The variance of the position in such systems grows over time, even if the mean particle position goes to zero for long observation times or many noise realizations. For this reason, the differencing method has to be applied in order to obtain the correct numerical results.
@@ Appendix C Extension to a system with harmonically coupled probe particle
Here we consider a one-dimensional system of a Brownian particle subject to time-delayed feedback, which is coupled to a passive probe particle via an attractive harmonic potential. We refer to the feedback-driven particle using the superscript $\mathrm{fb}$ and to the tracer particle using the superscript $\mathrm{pr}$. The dynamics of the feedback-driven particle is given by an overdamped Langevin equation. The tracer particle is coupled to the feedback driven particle via the position $x^{\mathrm{fb}}$, which enters the harmonic coupling potential. The coupled equations of motion then read
@@
resulting in the force $F_{\mathrm{c}}=-\nabla_{x_{\mathrm{pr}}}U$ and analogously $-\nabla_{x_{\mathrm{fb}}}U=-F_{\mathrm{c}}$ which acts on the coupled probe particle. Furthermore, $\xi_{\mathrm{fb}}$ and $\xi_{\mathrm{pr}}$ represent a one-dimensional Gaussian white noise, respectively, with zero mean and correlation function $\langle\xi_{\mathrm{fb}/\mathrm{pr}}(t)\xi_{\mathrm{fb}/\mathrm{pr}}(t^{\prime})\rangle=2\gamma_{\mathrm{fb}/\mathrm{pr}}k_{B}T\delta(t-t^{\prime})$. Then, the Fourier transform of the dynamics reads as follows.
@@
$-k[\tilde{x}_{\mathrm{fb}}(\omega)-\tilde{x}_{\mathrm{pr}}(\omega)]+\tilde{\xi}(\omega),$ (36)
@@
$\tilde{x}_{\mathrm{fb}}$ $=\frac{\frac{k\tilde{\xi}_{\mathrm{pr}}}{\gamma i\omega+k}+\tilde{\xi}}{\gamma i\omega+k_{a}-k_{b}e^{-i\omega\tau}+k-\frac{k^{2}}{\gamma i\omega+k}},$ (38)
@@
The analytical expression for the PSD then follows from the definition as for the simple linear system. From Eqs. (38) and (39) we can already see that the characteristic oscillations, and, thus, the signature of time delay extend to the probe particle. Therefore and in principle, by observing the probe particle alone, inference of time delay is possible with a method similar to that proposed in Sec. V the main text.

### v2

@@ B.1 Numerical solution of the overdamped delay Langevin equation
where $N_{\tau}=\tau/\Delta t$ is the number of time steps within one delay time $\tau$, for the discretization time step $\Delta t$. $\eta_{n}$ is a uncorrelated Gaussian random number with zero mean and unit variance, i.e., $\langle\eta_{n}\eta_{m}\rangle=0$ for $n\neq m$. Equation (29) is solved provided an initial history for the time interval $\tau,0]$ via the discretized history function $\Phi(t_{n})=\Phi_{t_{n}}$, that is, the trajectory in the discretized time interval $\tau,0]$, has to be known. For all stochastic simulations in this study, we used the trajectories of free Brownian particles as the initial history. Furthermore, we used time steps of $\Delta t=10^{-4}\tau_{B}$ (much smaller than the smallest delay time considered) to ensure sufficient numerical precision in all the cases considered.
B.2 The Differencing method
@@
Figure B1 presents exemplary data for systems with attractive and repulsive harmonic time-delayed feedback which both behave diffusively, especially in the long time limit. The cyan and red data correspond to PSD computed by directly applying the Fourier transform to the nonstationary positional data.
Figure B1: Comparison of position based and differencing based computation of $\mathrm{PSD}$ and $\omega^{2}\mathrm{PSD}$ for a system with linear feedback force [Eq. (2)]. Panels (a) and (c): Marginally stable system with attractive harmonic feedback; panels (b) and (d) are for repulsive harmonic feedback below the deterministic threshold, i.e., $|k_{a}|=|k_{b}|<\gamma/\tau$.
We observe a shift compared to the analytical expectation that is hardly visible in the PSD but becomes apparent in $\omega^{2}$PSD. While when averaging over a small number of noise realizations, this shift is lost in noise, when averaging over a larger number of noise realizations, the shift becomes obvious.
The reason behind this behavior can be understood when considering that we are taking the Fourier transform of a (within the observation period) not strictly ergodic and certainly non-stationary quantity. More specifically, the time series is non-stationary since the variance of the position grows over time. Changing the temporal resolution or the length of the trajectory only has a small effect on the observed shift. The observed shift can intuitively be understood as follows: In a stationary system (assuming zero mean and constant variance), the position will (more or less) fluctuate around zero, given that the observation window is long enough. However, if the process is not stationary in the sense that the variance grows as a function of time, in some realizations, the position will be positive (or negative) for a significant part of the observation time. When taking the Fourier transform of such a trajectory, it will be somewhat biased. If we took the average over noise realizations now, we could expect this bias to average out over different realizations. However, following the definition of the PSD (7) we compute the square modulus of the Fourier transform of the position before applying the noise average. For a stationary process, this does not matter. However, in our case, the bias can not be averaged out but rather accumulates and depends on the variance of the process and how much of the observation period will be uninterruptedly spent on one side of the mean position. We therefore have to approach the problem differently and make the time series stationary before even computing the Fourier transform in order to obtain results that match our analytical expectation.
@@
where the operator $\mathcal{F}$ represents the Fourier transform. Intuitively, this looks like the Fourier transform over a temporal derivative (which we already exploited in the analytical considerations) or essentially (for small $\Delta t$) the Fourier transform of the right-hand side of the discretized Langevin equation. Applying this method, we can obtain a numerical result that matches exactly the analytical expectation, as shown by the dark blue curves in Fig. B1.
The differencing method was used for all results discussed in Sec. IV of the main text.
B.3 Extraction of frequencies corresponding to minima and maxima of the PSD
In the following, we describe a procedure that allows us to extract the delay time from simulation data, which was used in Sec. IV.4 of the main text. It is based on the analytical results obtained for the linear system, but can also be applied to systems that depend on $x(t)$ and/or $x(t-\tau)$ in a nonlinear fashion. This procedure relies on the spacing of the positions of maxima and minima in the quantity $\omega^{2}\mathrm{PSD}$, and is applicable as long as the signal-to-noise ratio in the system allows for reliable determination and extraction of the position of these minima and maxima. As such, the larger the number of available noise realizations, the better the inference results. For example, Fig. B1 illustrates the level of noise that has to be dealt with in this kind of system in the case of a marginally stable time delay system. In order to obtain a good polynomial fit over multiple peaks, we need to average over a sufficient number of noise realizations; for the figures in the main text [Figs. 6a-c] 400 noise realizations were used.
While, in principle, the positions of the minima can be obtained manually, it is more reliable and convenient to use a spline interpolation for a subsequent semi-automatic extraction of the positions of minima and maxima. It is semi-automatic in the sense that one has to determine for how many minima and maxima a reliable fit can be obtained. Specifically, we use quartic splines for fitting the numerical data, which allows us to extract the roots of its derivative automatically. As the data are noisy and somewhat irregular, special care has to be taken when adjusting the smoothing parameter in the spline interpolation, such that the interpolation faithfully represents the location of the minima and maxima in the numerical data. Furthermore, limiting the range of the interpolation can significantly improve the results since at low frequencies additional oscillations with different oscillation period are observed in some cases and at high frequencies the oscillations are lost in noise.
@@ B.3 Extraction of frequencies corresponding to minima and maxima of the PSD
    Compute the roots $\omega_{n}$ of the derivative of the polynomial corresponding to the minima and maxima in $\omega^{2}{\rm PSD}$ (built-in function of spline interpolation in scipy).
@@
$2\pi/\tau_{\mathrm{inferred}}=\frac{\sum_{n=1}^{N_{\mathrm{ext}}-2}\left(\omega_{n+2}-\omega_{n}\right)}{(N_{\mathrm{ext}}-2)}\ ,$ (31)
where $\tau_{\mathrm{inferred}}$ is the inferred delay time, $N_{\mathrm{ext}}$ is the number of minima and maxima that can be reliably identified, and we averaged the distances between all available pairs of consecutive minima and maxima.
B.4 Deep learning: data preparation
@@
The data are prepared in this way for training the neural network as well as for inference using the trained network (see Sec. VI).
B.5 Deep learning: network architecture and hyperparameter tuning
For the classification of trajectory data we use a convolutional deep neural network (see e.g., [55, 56]) implemented using keras [98] based on tensorflow [99]. The neural network (see Sec. V.2, especially Fig. 9 for a simplified schematic depiction) consists of an input layer (the number of nodes equals the number of relevant time steps), hidden layers, and an output layer (the number of nodes equals the number of classes, i.e., delay times the network is trained on). For hidden layers, considerations such as complexity, regularization, and available resources play a crucial role. Here, we focus on a three convolutional layers and pooling layers to limit complexity. We use a hyperparameter tuning scheme in order to achieve the desired accuracy for classification by the network. This means that over how many nodes pooling is applied in a pooling layer, filter and kernel size in convolutional layers, the optimization method and the activation function are chosen automatically. Before the final hidden layer, the network is flattened. The final hidden layer is then a densely connected layer followed by the densely connected output layer. The number of nodes of the output layer is equal to the number of different delay times on which the model is trained. A softmax activation function is used to obtain a classification in terms of the probability that the classified trajectory is based on a certain delay time. The network is then constructed as shown in Table 2.
@@ B.5 Deep learning: network architecture and hyperparameter tuning
Table 2: Network structure by layers; hyper-parameters, i.e., over how many nodes pooling is applied in a pooling layer, filter and kernel size in convolutional layers, the optimization method and the activation function are optimized by hyperparameter tuning. For the input data preparation, see Appendix B.4.
@@
As the loss function, which is minimized during training, the sparse categorical cross-entropy was used. Figure B2 presents exemplary accuracy [Fig. B2(a)] and loss [Fig. B2(b)] as a function of training epochs for the training of the neural network on several nonlinear feedback forces [Eqs. (11)-(13)] simultaneously. We observe good convergence already at as little as 20 training epochs.
@@
Figure B2: Training the convolutional neural network on nonlinear feedback forces [Eqs. (11)–(13)]: Panel (a) shows the classification accuracy, and panel (b) shows the training loss (computed via sparse categorical crossentropy), both plotted as functions of training epochs for the training and validation datasets.
B.6 Characteristic response of a nonlinear feedback force system
@@
Figure B3 presents exemplary data for the response of systems with nonlinear feedback forces. In all cases, a complex but still characteristic response is observed. As for linear systems, a change in dynamical behavior is observed one delay time after switching on the perturbation as well as one delay time after switching off the perturbation.
@@
Figure B3: Characteristic response of a nonlinear system with time delay to an external perturbation (see also Sec.V.1). Panel (a): Response of the $\tanh$ feedback force [Eq.(3)] to a constant external force. Panel (b): The response of a system based on Eq. (13) to the same type of perturbation. Vertical dotted lines indicate the onset and termination of the perturbation, while the vertical dashed line marks the end of the observation window.
B.7 Characteristic response of a stochastic system with nonlinear feedback force
Figure B4 presents exemplary data for the response of stochastic systems (11)-(13) subject to nonlinear and delayed forces with corresponding inference results. In contrast to the linear system (for comparison see Fig. 8), the response of a nonlinear system can be more complex, as indicated by the solid lines representing the deterministic response in Fig. B4. In the stochastic data, the changes in dynamical behavior that constitutes the characteristic response is obfuscated by thermal noise. Still, the neural network can infer time delay with high accuracy for delay times the network was trained on for trajectories prepared as discussed in Appendix B.4. For delay times in-between two subsequent $\tau_{\mathrm{node}}$, the inference results tend to be less accurate.
Figure B4: Characteristic response of a stochastic nonlinear system with time delay to an external perturbation (see also Sec. V.1) and corresponding inference results (right panels). Trajectories prepared as discussed in Appendix B.4. Solid line: corresponding deterministic response. Dashed horizontal line: true delay time. Panels a)-c): $\tau=0.15\tau_{B}$; a) Eq. (11) at $a=13k_{\mathrm{B}}T/\sigma^{4}$, $b=3k_{\mathrm{B}}T/\sigma^{2}$, $c=11k_{\mathrm{B}}T/\sigma^{2}$; b) Eq. (12) at $a=13k_{\mathrm{B}}T/\sigma^{4}$, $b=18k_{\mathrm{B}}T/\sigma^{2}$, $c=17k_{\mathrm{B}}T/\sigma^{3}$; c) Eq. (13) at $a=7k_{\mathrm{B}}T/\sigma$, $b=10k_{\mathrm{B}}T/\sigma$, $c=17k_{\mathrm{B}}T/\sigma$. Panels d)-f): Trajectories based on Eq. (13) at d) $\tau=0.0125\tau_{B}$, $a=7k_{\mathrm{B}}T/\sigma$, $b=10k_{\mathrm{B}}T/\sigma$, $c=17k_{\mathrm{B}}T/\sigma$; e) $\tau=0.075\tau_{B}$, $a=13k_{\mathrm{B}}T/\sigma$, $b=3k_{\mathrm{B}}T/\sigma$, $c=11k_{\mathrm{B}}T/\sigma$; f) $\tau=0.15\tau_{B}$, $a=9k_{\mathrm{B}}T/\sigma$, $b=3k_{\mathrm{B}}T/\sigma$, $c=7k_{\mathrm{B}}T/\sigma$.
@@ Appendix C Extension to a system with harmonically coupled probe particle
Here we consider a one-dimensional system of a Brownian particle subject to time-delayed feedback, which is coupled to a passive probe particle via an attractive harmonic potential. We refer to the feedback-driven particle using the superscript $\mathrm{fb}$ and to the probe particle using the superscript $\mathrm{pr}$. The dynamics of the feedback-driven particle is given by an overdamped Langevin equation. The tracer particle is coupled to the feedback driven particle via the position $x_{\mathrm{fb}}$, which enters the harmonic coupling potential. The coupled equations of motion then read
@@
resulting in the force $F_{\mathrm{c}}=-\nabla_{x_{\mathrm{pr}}}U$ and analogously $\nabla_{x_{\mathrm{fb}}}U=-F_{\mathrm{c}}$ which acts on the coupled probe particle. Furthermore, $\xi_{\mathrm{fb}}$ and $\xi_{\mathrm{pr}}$ represent a one-dimensional Gaussian white noise, respectively, with zero mean and correlation function $\langle\xi_{\mathrm{fb}/\mathrm{pr}}(t)\xi_{\mathrm{fb}/\mathrm{pr}}(t^{\prime})\rangle=2\gamma_{\mathrm{fb}/\mathrm{pr}}k_{B}T\delta(t-t^{\prime})$. Then, the Fourier transform of the dynamics reads as follows.
@@
$-k[\tilde{x}_{\mathrm{fb}}(\omega)-\tilde{x}_{\mathrm{pr}}(\omega)]+\tilde{\xi}_{\mathrm{fb}}(\omega),$ (35)
@@
$\tilde{x}_{\mathrm{fb}}$ $=\frac{\frac{k\tilde{\xi}_{\mathrm{pr}}}{\gamma i\omega+k}+\tilde{\xi}_{\mathrm{fb}}}{\gamma i\omega+k_{a}-k_{b}e^{-i\omega\tau}+k-\frac{k^{2}}{\gamma i\omega+k}},$ (37)
@@
The analytical expression for the PSD then follows from the definition as for the simple linear system. From Eqs. (37) and (38) we can already see that the characteristic oscillations, and, thus, the signature of time delay extend to the probe particle. Therefore and in principle, by observing the probe particle alone, inference of time delay is possible with a method similar to that proposed in Sec. V of the main text.
Appendix D Systems with exponentially correlated noise – numerical results
In this section, we demonstrate that our methods of inferring the time delay are not limited to white noise systems but can be applied or extended to systems with colored noise. In other words, the signatures of time delay remain intact even in colored-noise systems.
To this end, we first introduce exponentially correlated colored noise and subsequently present exemplary data for the (power spectral density) PSD-based inference approach as well as the neural network-based approach for non-linear systems. Specifically, we consider the nonlinear systems in in Eqs. (11)-(13) and replace the Gaussian white noise (present on their right-hand side) by an additive colored noise [103]
$\dot{\zeta}(t)=-\frac{1}{t_{\mathrm{corr}}}\zeta(t)+\sqrt{\frac{2\gamma k_{\mathrm{B}}T}{t_{\mathrm{corr}}^{2}}}~\eta(t)\ ,$ (39)
where $t_{\mathrm{corr}}$ is the correlation time and $\eta(t)$ is a white noise (delta-correlated) with zero mean and unit variance.
It turns out that in the long-time limit (i.e., the stationary state), the colored noise $\zeta(t)$ is exponentially correlated:
$\langle\zeta(t)\zeta(t^{\prime})\rangle_{\rm ss}=\frac{\gamma k_{\mathrm{B}}T}{t_{\mathrm{corr}}}e^{-|t-t^{\prime}|/t_{\rm corr}}\ .$ (40)
Before proceeding further, we would like to highlight its two limiting behaviors. 1) In the limit $t_{\mathrm{corr}}\to 0$, for $t=t^{\prime}$, the covariance (40) diverges; otherwise, the correlation goes to 0 for $t\neq t^{\prime}$. Further, integrating the right-hand side of Eq. (40) from $t\in(-\infty,+\infty)$ gives $2\gamma k_{\rm B}T$. Therefore, this correlation is a delta function in this particular limit (with a prefactor $2\gamma k_{\mathrm{B}}T$). 2) In the limit $t_{\mathrm{corr}}\to\infty$, $\dot{\zeta}\to 0$, i.e., $\zeta$ becomes a quenched random Gaussian variable with zero mean and variance $\gamma k_{\mathrm{B}}T/(t_{\rm corr})\to 0$ (as $t_{\rm corr}\to\infty)$.
Figure D5 shows the PSD and the extracted positions of minima and maxima. Table 3 presents the corresponding inference results for one delay time at different values of the correlation time $t_{\mathrm{corr}}$. Interestingly, while the behavior of the PSD as a function of the frequency changes as the colored noise is added to the system and the correlation time $t_{\mathrm{corr}}$ is increased, the characteristic signature of time delay, i.e., oscillations with a delay-specific oscillation period, remains.
Figure D5: Numerically simulated PSD (red) as a function of $\omega$ for non-linear systems (12) with white noise replaced by colored noise. In each panel: Polynomial fit (blue) using the quartic splines and automatically extracted positions (cyan circles) of maxima and minima using the roots of the derivative of the polynomial (see Appendix B.3). See Table 3 for the corresponding inference results. Colored noise at correlation times: a) $t_{\mathrm{corr}}=0.01\tau$, b) $t_{\mathrm{corr}}=0.1\tau$, c) $t_{\mathrm{corr}}=1\tau$ and d) $t_{\mathrm{corr}}=10\tau$, other parameters as in Fig. 6 and Table 1.
$t_{\mathrm{corr}}/\tau$ $\tau/\tau_{\rm B}$ $\tau_{\mathrm{inferred}}/\tau_{\rm B}$ $\sigma_{\tau}/\tau_{\rm B}$
$0.01$ 0.15 $0.151$ $\pm 0.003$
$0.1$ 0.15 $0.152$ $\pm 0.007$
11 0.15 $0.153$ $\pm 0.009$
$10$ 0.15 $0.152$ $\pm 0.010$
Table 3: Inference of the delay times corresponding to system in Eq. (12) based on the PSD-data shown in Fig. D5.
For the neural network approach, Fig. D6 shows data for inference results of the modified nonlinear systems in Eqs. (11)-(13) at two different correlation times using a network trained solely on white noise systems. In the white noise limit, i.e., $t_{\mathrm{corr}}\ll\tau$, the inference results resemble those obtained for white noise systems. However, when the correlation time is of the order of the delay time or longer, the resulting trajectories look significantly different than before. Thus, as we expect, a neural network trained on systems with only white noise may not be able to accurately infer the time delay for systems with longer correlation times. Therefore, one requires to train the neural network with different correlation times (including different combinations of other parameters) to have better accuracy when inferring the time delay.
Figure D6: Inference results for a network trained on non-linear systems [Eq. (11)-(13)] with white noise, for trajectories from colored noise systems based on [Eq. (11)-(13)] at correlation times $t_{\mathrm{corr}}=0.001\tau$ [panels a) and b)] and $t_{\mathrm{corr}}=0.1\tau$ [panels c) and d)]. Panels (a) and (c): Symbols (color intensity increasing with $\tau_{\rm node}$): The probability of inferring a delay time at an exit node corresponding to $\tau_{\mathrm{node}}$. Average over inference results for ten different trajectories (prepared as discussed in Appendix B.4) for each parameter combination. Each data point is an average over 27 different combinations of the system parameters. Error bars indicate one standard error of the mean. Vertical dotted lines indicate $\tau_{\rm node}$. Panels (b) and (d): Same data (but log-scaled $y$-axis) as in panel (a) and (c) represented by scaling and shifting the data to reflect the deviation from the output node $\tau_{\rm node}$. Grey Gaussian curve centered around zero as well as horizontal grey line (indicating probability one) act as guides for the eye.